# Improving the Reasoning of Multi-Image Grounding in MLLMs via Reinforcement Learning

### v1

@@ 5 Conclusion
In this paper, we aim to enhance the reasoning and understanding capabilities of MLLMs for real-world multi-image grounding applications. To this end, we propose a post-training strategy incorporating cold-start initialization to establish correct reasoning pathways, complemented by rule-based reinforcement learning to further strengthen the model’s reasoning abilities. Experimental results demonstrate that our approach significantly outperforms both the base model and the baseline SFT method across multiple multi-image grounding and understanding benchmarks. We hope this work will stimulate further research in the multi-image reasoning and grounding community.

### v2

@@ 5 Conclusion
In this paper, we aim to enhance the reasoning and understanding capabilities of MLLMs for real-world multi-image grounding applications. To this end, we propose a post-training strategy incorporating cold-start initialization to establish correct reasoning pathways, complemented by rule-based reinforcement learning to further strengthen the model’s reasoning abilities. Experimental results demonstrate that our approach significantly outperforms both the base model and the baseline SFT method across multiple multi-image grounding and understanding benchmarks. We hope that this work will inspire further research in the multi-image reasoning and grounding community.